OpenStudy (anonymous):

there are three terms of arithmetic sequence a1=decrease by 5 a2=decrease by 1 a3=twice the original term find the a3..???

OpenStudy (anonymous):

LoL Derek ! XD

hartnn (hartnn):

is the common difference also given ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its 11

OpenStudy (anonymous):

Given yung Common Difference ! 11 :P

hartnn (hartnn):

common difference = A2-A1 = A3-A2 = 11

OpenStudy (anonymous):

the common difference is 11, sorry. xD

OpenStudy (agent0smith):

This is written confusingly... a1=decrease by 5 a2=decrease by 1 a3=twice the original term decrease what by 5??

hartnn (hartnn):

A3-A2 = 11 2n - (n-1) =11 can you find 'n' from here ??

OpenStudy (anonymous):

the first term is decreased by5

hartnn (hartnn):

true! wording are really confusing and question is really weird!

OpenStudy (agent0smith):

The first term is decreased by 5 to make what... ? the second term? itself? It doesn't make a lot of sense.

OpenStudy (anonymous):

yup it really doesn't

hartnn (hartnn):

@agent0smith look at this also http://openstudy.com/study#/updates/51138785e4b0e554778acff1 better notations

OpenStudy (anonymous):

yup it really doesn't make any sense

OpenStudy (agent0smith):

That one makes a lot more sense. Lets use that question.

hartnn (hartnn):

unfortunately no... thatsweird too... A2-A1 is not 11 actually....

OpenStudy (anonymous):

hahahaha. xD that's my classmate. xD we have the same question

hartnn (hartnn):

lol .ok :)

OpenStudy (anonymous):

Yes !

hartnn (hartnn):

and is ploeng in love with derek ?? O.o

OpenStudy (anonymous):

hahaha. xD NO!!!!!!!!!!!!!!!

OpenStudy (anonymous):

nope :))

hartnn (hartnn):

okk. just wondering...

OpenStudy (agent0smith):

The other question seems broken too...

OpenStudy (anonymous):

Yah ! I think so .

hartnn (hartnn):

thats what i said...

OpenStudy (anonymous):

2n-n-1=11 n=12 substitute it gave me 7,11 and 24

OpenStudy (anonymous):

doesn't match with the common difference

hartnn (hartnn):

ok...i guess, 11 as common difference(d) should naot have been given. in that case d = A2 - A1 = 4 then A3-A2 = 4 and find 'n'.

hartnn (hartnn):

n=3 will give u a3 = 2n = 2*3 =6